Standing waves are established in a string. The frequency of the wave is 512 Hz and the distance between two successive nodes is 8 cm. Calculate the velocity of the wave in the string.


1
Expert's answer
2021-05-10T02:42:31-0400

velocity of wave = frequency "\times" wavelength

wavelength = 8 cm or 0.08 m

Frequency= 512 Hz


"\nu=\digamma\lambda"

the distance between two nodes represents half wavelength = 0.5 "\lambda"


wavelength = "2\times 0.08 = 0.16 m"


"\nu= 512 \times 0.16 = 81.92 m/s"